OpenStudy (anonymous):

Assume that each born child is equally likely to be a boy or a girl. If a family has two children, what is the conditional probability that both are girls given that (i) the youngest is a girl, (ii) at least one is a girl?

OpenStudy (anonymous):

this is the kind of problem where you can write out all possibilities since there aren't that many: having 2 kids. there are only 4 possible outcomes: boy, boy girl, girl boy, girl girl, boy [i] youngest is a girl: 2 of the 4 outcomes have girl second. so the probability is 2/4 = 1/2 = 50% [ii] 3 of the 4 possibilities satisfy at least one being a girl. answer is 3/4 or 75%

OpenStudy (anonymous):

girl 2nd means she is younger? like b,g means boy is older and girl is younger?